i need help!!! Im stuck on this word problem.

Three Pairs of jeans and six shirts caost $104.25. The cost of 4 jeans and 5 shirts is $112.15. Find the cost of each pair of jeans and each shirt.

: i need help!!! Im stuck on this word problem.

Three Pairs of jeans and six shirts caost $104.25. The cost of 4 jeans and 5 shirts is $112.15. Find the cost of each pair of jeans and each shirt.


Answer by Alan3354(1418)

Three Pairs of jeans and six shirts caost $104.25. The cost of 4 jeans and 5 shirts is $112.15. Find the cost of each pair of jeans and each shirt.
-------------------
Start by writing down what it says.
3J + 6S = 10425 eqn 1 from sentence 1
4J + 5S = 11215 eqn 2 from sentence 2
Now you have 2 eqns in 2 unknowns. There are several ways to solve them.
I'll use elimination
Multiply the 2 eqns to get the same number of one of the variables.
12J + 24S = 41700 eqn 1 x 4
12J + 15S = 33645 eqn 2 x 3
Subtract
0J + 9S = 8055
S = 895
------------Sub for S in eqn 1
3J + 6*895 = 10425
3J + 5370 = 10425
3J = 5055
J = 1685
So the shirts cost $8.95 and the jeans are $16.85